# Robot Localization aided by Quantum Algorithms

### v1

@@ 2.2.1 Grover’s quantum searching algorithm
Grover’s quantum searching algorithm is usually described (Grover (1996)) in terms of the iteration of a compound operator Q of the form (Jozsa (1999))
@@
On a starting state |$\ket{\psi_{0}}=H\ket{0}$. Here $H$ is the Walsh-Hadamard transform and $I_{0},I_{x_{0}}$ are suitable inversion operators. The operator $-HI_{0}H$ was originally called a “diffusion” operator (Grover (1996)) and later interpreted as “inversion in the average” (Grover (1997)).
@@
Here the input register |$\ket{x}$ consists of $n$ qubits as $x$ ranges over all $n$ bit strings and the output register |$\ket{y}$ consists of a single qubit with $y=0$ or 1. The symbol $\oplus$ denoted addition modulo 2 ("modulo" is the remainder of a division operation between two numbers).
The assumption that the database was unstructured is formalised here as the standard oracle idealisation that we have no access to the internal workings of $U_{f}$ – it operates as a “black box” on the input and output registers.
In this formulation there is no problem with the access to $f(x)$ for any of the exponentially many $x$ values and indeed we may also readily query the oracle with a superposition of input values.
Instead of using $U_{f}$, an equivalent operation is used. Such operation is denoted $I_{x_{0}}$ on $n$ qubits. It is defined by
@@
Our searching problem becomes the following: we are given a black box which computes $I_{x_{0}}$ for some $n$ bit string $x_{0}$ and we want to determine the value of $x_{0}$.
Given the black box $I_{x_{0}}$ how can we identify x0? Grover’s Algorithm 2 states that, in order to locate (in an unordered list) the elements that we are looking for, we should follow the algorithm:
@@
Figure 6 shows the main steps of the Grover algorithm:
@@
With this algorithm at hand, we can try to check if it suits our localization problem. If we take Figure 4 and write the numbers in binary form, we obtain Figure 7
@@
Figure 7: : The robot localization should be done trying to match the perception costmap (left), with the previously generated costmap (right)
We can try to check if this algorithm could be used to find the position of a certain substring (obtained by the sensors of the robot) in a longer string (where the environment map is encoded). For example, we can use Qiskit to implement Grover’s algorithm for such detection. They key here is to be able to define an Oracle that is based in a function that achieves f($x_{0}$)=1, and then apply Algorithm 2.
If we use "10110001" as a certain row where we look for the pattern "01", and use Grover’s algorithm we obtain the results shown in Figure 8:
@@ 3 Proposed approach
Grover’s algorithm offers a powerful quantum computing approach to searching problems, and it can be adapted for robot localization within a known map by dramatically reducing the time complexity associated with finding the correct position. In a typical robot localization problem, the goal is to identify the robot’s position within a predefined environment. This environment can be represented as a map with a finite number of discrete locations, each representing a possible position where the robot might be.
@@
The algorithm has been implemented using Qiskit (ibm (2024a)), in a way that associates each candidate location on the map with a quantum state, where the correct state (the robot’s actual location) is the one marked by an “oracle” function. This oracle has been programmed to recognize and mark the actual location by matching the candidate state to the real-time data collected from the robot’s sensors. The oracle then enhances the amplitude of the state representing the correct location. By repeatedly applying the Grover operator, the probability amplitude of the correct position is amplified with each iteration, making it more likely to be measured upon observation.
In practical terms, (if real quantum computing achieves enough capacity) this quantum-enhanced localization approach could be particularly beneficial in environments with high location ambiguity or large search spaces, such as a warehouse with many aisles or complex navigation zones. In principle, leveraging quantum computing’s ability to quickly process great amounts of data (exponentially faster than classical computers), robots and autonomous vehicles can quickly determine their precise location and orientation, even in situations where GPS signals are weak or unavailable.
@@
That is why in this paper we propose using Grover’s quantum search oracle for performing a double localization in a 2D map: one for the horizontal axis and another for the vertical axis, as shown in Figure 9
Figure 9: : Block diagram of proposed approach
@@
We left aside other current approaches for pattern matching, like the use of Parametric Probabilistic Quantum Memory (Sousa et al. (2020)) (PQM, a data structure that computes the Hamming distance from a binary input to all binary patterns stored in superposition on the memory), as they require the substring to be searched to be the same length as the strings in the stored data.
@@ 3.1 Algorithm
Figure 10: : The perception costmap (left) has 3 horizontal elements, and 3 vertical elements. The map (the previously generated costmap, in the right) is a 4x4 grid of cells (each cell has two qubits)
@@
Figure 11: : Considering (0,0) as the upper left cell, for coordinates (1,1) if the map’s cell values (in black) match the values for the horizontal (blue) and vertical (red), the Oracle’s output should be ’1’, and ’0’ otherwise. In other words: the algorithm should look for [11,10,11] horizontally, and [11,10,10] in vertical.
The code that has been generated in this paper (Antero (2024)) can be configured for any map size, any number of qubits per cell, and any row or column pattern. In this case (with a 4x4 map with 2 qubits per cell, and horizontal and vertical patterns with size 3 (using 2 qubits per cell), the Oracle is automatically generated. Such oracle (in this case) has a possible search space of: [(1,1), (1,2), (2,1), (2,2)], and would return ’1’ in the output qubit for position: (1,1).
@@ 4 Experimental Setup and Validation
In the preceding sections, we introduced a novel application of quantum computing to the problem of robot localization, proposing a quantum algorithm designed to efficiently solve the localization problem. To validate the feasibility and effectiveness of this quantum approach, this section presents a comprehensive experimental setup and validation framework. We outline the methodology employed to evaluate the performance of our quantum algorithm in estimating a robot’s position and orientation within a given environment, and present the results obtained from a series of simulations and experiments. These experiments aim to assess the accuracy, robustness, and scalability of our quantum algorithm, and demonstrate its potential advantages over classical methods for robot localization. The remainder of this section details the experimental design, performance metrics, and results, offering a thorough analysis of the proposed quantum approach to robot localization and its potential applications in robotics and autonomous systems.
@@
The first module, is an auxiliary configuration tool developed using Pygame (pyg (2024)) (a Python library designed for creating video games and multimedia applications). This tool enables users to define and customize the parameters of the localization problem, including the 2D map, robot properties, and algorithmic settings. The output of the tool is a standardized input file that serves as the basis for the second module.
@@
  • Create a map with a size $nxn$, being $n\geq 2$
@@
Once launched the application shows a simple user interface where, using arrow keys, it is possible to move the robot to any cell in the map.
@@
Using this user interface, different setups (described in the next subsections) have been prepared and tested. These setups cover different situations: from small grids that were simulated locally in a PC, to more complex grids that had to be simulated in real quantum computers.
On the other hand, in the scope of this paper we also have implemented the algorithm described in 3. This main module, built using the Qiskit quantum development environment, implements the quantum algorithm for robot localization in a 2D map.
@@
By decoupling the problem configuration from the quantum algorithm implementation, our framework provides a versatile and user-friendly environment for testing and evaluating the performance of the proposed quantum algorithm for robot localization.
In any case, it should be noted that quantum computing, as a technology, is still in its infancy. It is clear that nowadays different engineering challenges in the real quantum computer disable (in a certain sense) any advantage that the proposed algorithm may present. In another words: notwithstanding the theoretical validity of the proposed quantum approach, the practical implementation of the corresponding quantum circuit is subject to the engineering limitations of current quantum computing technology.
@@
In real quantum computers we currently face four main challenges:
  • Qubit state stability: the stability of a qubit’s state typically lasts only a few hundred microseconds, which limits the duration of computations.
@@
  • Limited qubit availability: the number of physical qubits available in these devices remains constrained, posing a significant bottleneck for scalability.
Limited qubit availability impacts on the complexity of the algorithms we can use, but other three issues directly impact on the error probability of the final result. As indicated in (Johnstun and Huele (2021)), it is important to understand the effect of these issues and have some estimation of the final error.
In this paper we use data provided by the manufacturer of the real quantum computer (IBM in our case), to calculate the total error probability expected during the run of a given quantum circuit. Using the methodology (and code) proposed in (Aseginolaza et al. (2024)), for each setup we estimate how far are we from the limits of current quantum computers (for using their code in this paper, the code has been updated to the latest version of Qiskit).
In the next subsections, we describe the experimental setups and results obtained using this software framework, which demonstrate the effectiveness of our quantum approach to robot localization (and the limits of current physical quantum computers).
4.1 Very small grid, with simple pattern match
As a first test for validation, a very simple grid has been configured: the grid uses a single qubit per cell (the calculated costmap in each cell has been configured to use a single qubit), and a 2x2 qubit grid has been defined.
The Pygame based interface has been used to configure such grid, and to define the robot’s position in a certain cell in the grid.
Using the Pygame interface makes it possible to easily simulate how a Grover-based quantum localization algorithm that would search for a target within the grid, moving through cells and marking positions.
The visual grid provides an intuitive way to observe the robot’s position and how its sensors can perceive its surrounding, making it useful for testing and experimentation.
The robot has been positioned in an arbitrary location, as seen in Figure 13.
Map qubits: $\begin{bmatrix}0&0\\ 0&1\\ \end{bmatrix}\$
@@ 4.1 Very small grid, with simple pattern match
Figure 13: : Simple setup with a 2x2 grid
The proposed double Grover search with a pattern $[1,0]$ in a 2x2, has to deal with a search space of 4qubits.
For this setup our algorithm requires to use:
  • 2 qubits for the search space (1 for encoding the horizontal position, and 1 for the vertical one)
  • The Oracle has to deal with 4 possible positions (2x2), with an expectation of a single correct solution (N=4, M=1)
@@
  • The total number of required qubits is 10
In order to validate the effectiveness of the algorithm, the Qiskit code has been tested:
  • Simulating the circuits locally in a computer using the ’qasm’ simulator
  • Using the IonQ Quantum Cloud (ion (2024)), where it’s possible to simulate circuits as if they were running on IonQ hardware, using a special noise model simulation.
  • Using the QuantumInspire Cloud simulator (qua (2024)), where it’s possible to simulate circuits using their simulators.
  • Using a real quantum computer from IBM simulator (ibm (2024b)). Using IBM’s SDK it’s possible to send a quantum circuit to a real IBM quantum computer.
Grover’s algorithm clearly identified the correct solution: row 1 and column 1, which is encoded as position ’3’ or ’11’ in binary (count starts from zero). The correct solution’s probability was clearly amplified in the simulated environments.
It should be noted that in the real quantum computer, although the correct solution was found, the amplification is not so evident. Current quantum computers operate in the realm of quantum mechanics, governed by principles of superposition, entanglement, and interference. However, as it was previously mentioned, these systems are prone to noise (which refers to random errors that occur during computation), causing qubits to lose their quantum properties and behave classically. As a result, actual computations often deviate from idealized simulations, leading to discrepancies between theoretical predictions and experimental results.
The obtained histograms are shown in Figure 14:
(a) QasmSimulator
(b) QuantumInspire
(c) IonQ simulator
(d) Real IBM Quantum computer
Figure 14: : Histograms of the same quantum circuit obtained from simulated and real quantum computers
It is clear that the noise in the real quantum computer disables (in a certain sense) any advantage that the proposed algorithm may present. In the performed tests, although most of the times the results are correct, sometimes the algorithm tends to fail in the real computer. As previously mentioned, in order to measure how far is this setup from the current practical limits of quantum computers, using code by (Aseginolaza et al. (2024)) for this 2x2 grid we get:
  • Error due to noisy gates = 4.9%
  • Error due to estability = 99.91%
  • Measurement errors = 17.65%
  • Total estimated error probability $>99.96\%$
4.2 Small grid, with simple pattern match
To perform a more useful validation, it is necessary to perform a simulation that is computationally more demanding, so in this setup a 3x3 grid will be used. The setup contains a 2-qubit pattern search (both horizontally and vertically), and the calculated costmap in each cell has been configured to use a single qubit. Once again, the Pygame based interface has been used. The robot has been positioned in the central cell, as seen in Figure 15.
Map qubits: $\begin{bmatrix}0&0&0\\ 0&1&0\\ 0&0&0\\ \end{bmatrix}\$
Horizontal search qubits: $\begin{bmatrix}0,1\end{bmatrix}\\ \$
Vertical search qubits: $\begin{bmatrix}0,1\end{bmatrix}\$
Figure 15: : Simple setup with a 3x3 grid
The proposed double Grover search with a pattern $[0,1]$ in a 3x3, has to deal with a search space of $22x$ 2 (as the length of the pattern to be searched requires matching two cells). In order to evaluate the quantum circuit, the proposed algorithm requires to use 2 qubits, and the general requirements are:
  • 2 qubits for the search space (to encode positions 0 to 3)
  • The Oracle has to deal with 4 possible positions (2x2), with an expectation of a single correct solution (N=4, M=1)
  • The suggested number of repetitions is: $R=\pi/4\sqrt{N/M}=1$
  • The total number of required qubits is 18
  • Circuit depth is 35
@@ 4.2 Small grid, with simple pattern match
  • locally in a computer (using qasm)
  • using the IonQ cloud
  • using Bluequbit
@@
The real IBM was the "IBM Sheerbroke" machine, a 127qubit machine with a "2Q Error" rate (which is maeasured as the lowest two qubit error rate from all edges measured by isolated randomized benchmarking) of 2.66e-3, and a "2Q Error" rate (measured as the average two qubit error rate per layered gate for a 100-qubit chain) of 1.48e-2.
In this setup with a single obstacle in the middle, when the robot’s sensors measure ["0","1"] vertically, and ["0","1"] horizontally, the algorithm successfully interprets that this can only happen in the central point of the maze.
As the maze is a 3x3 grid, and the search items have a length of 2 cells, the search space is limited to a 2x2 space, which starts from the [1,1] coordinate cells. This means that in the search space, in order to get output coordinates the next translation has to be used (all coordinates are zero-based):
  • 00: translates to coordinates [1,1]
  • 10: translates to coordinates [1,2]
  • 11: translates to coordinates [2,1]
  • 01: translates to coordinates [2,2]
As shown in the results in 16, the algorithm proposes "00" as output, so the solution is [1,1], which is correct for the proposed map.
(a) QasmSimulator
(b) IonQ
(c) Bluequbit
(d) Real IBM Quantum computer
@@
Figure 16: : Histograms of the same quantum circuit obtained from simulated and real quantum computers
In order to validate the error probability due to use non-ideal real quantum computers, using the approach by (Aseginolaza et al. (2024)), for this 3x3 grid we get:
  • Error due to noisy gates $>50\%$
  • Error due to estability $>99.99\%$
  • Measurement errors $30\%$
  • Total estimated error probability $>99.99\%$
4.3 More complex grid, with a pattern match with two qubits

### v2

@@ 2.2.1 Grover’s quantum searching algorithm
Grover’s quantum searching algorithm is usually described ((Grover, 1996)) in terms of the iteration of a compound operator Q of the form ((Jozsa, 1999))
@@
On a starting state |$\ket{\psi_{0}}=H\ket{0}$. Here $H$ is the Walsh-Hadamard transform and $I_{0},I_{x_{0}}$ are suitable inversion operators. The operator $-HI_{0}H$ was originally called a “diffusion” operator ((Grover, 1996)) and later interpreted as “inversion in the average” ((Grover, 1997)).
@@
Here, the input register |$\ket{x}$ consists of $n$ qubits as $x$ ranges over all $n$ bit strings and the output register |$\ket{y}$ consists of a single qubit with $y=0$ or 1. The symbol $\oplus$ denoted the addition modulo 2 ("modulo" is the remainder of a division operation between two numbers).
The assumption that the database was unstructured is formalized here as the standard oracle idealization that we have no access to the internal workings of $U_{f}$ – it operates as a “black box” on the input and output registers.
In this formulation, there is no problem with the access to $f(x)$ for any of the exponentially many values of $x$, and indeed we may also readily query the oracle with a superposition of input values.
Instead of using $U_{f}$, an equivalent operation is used. This operation is denoted $I_{x_{0}}$ on $n$ qubits. It is defined by
@@
Our searching problem becomes the following: we are given a black box that computes $I_{x_{0}}$ for some $n$ bit string $x_{0}$ and we want to determine the value of $x_{0}$.
Given the black box $I_{x_{0}}$, Grover’s algorithm 2 states that in order to locate x0 (in an unordered list), we should follow the algorithm:
@@
Figure 6 shows the main steps of the Grover algorithm, with an initial preparation step, one block that contains the Oracle and the diffuser (which must be repeated), and a final measurement step.
@@
With this algorithm at hand, we can try to check if it fits our localization problem. If we take Figure 4 and write the numbers in binary form, we obtain Figure 7
@@
Figure 7: The robot localization should be done trying to match the perception costmap (left), with the previously generated costmap (right)
We can try to check if this algorithm could be used to find the position of a certain substring (obtained by the sensors of the robot) in a longer string (where the environment map is encoded). For example, we can use Qiskit to implement Grover’s algorithm for such detection. They key here is to be able to define an Oracle based on a function that achieves f($x_{0}$)=1, and then apply the Algorithm 2.
If we use "10110001" as a certain row where we look for the pattern "01", and use Grover’s algorithm, we obtain the results shown in Figure 8:
@@ 3 Proposed approach
Grover’s algorithm offers a powerful quantum computing approach to searching problems, and it can be adapted for robot localization within a known map dramatically reducing the time complexity associated with finding the correct position. In a typical robot localization problem, the goal is to identify the robot’s position within a predefined environment. This environment can be represented as a map with a finite number of discrete locations, each representing a possible position where the robot might be.
@@
The algorithm has been implemented using Qiskit ((ibm, 2024a)), in a way that associates each candidate location on the map with a quantum state, where the correct state (actual location of the robot) is the one marked by an ’oracle’ function. This oracle has been programmed to recognize and mark the actual location by matching the candidate state with the real-time data collected from the robot sensors. The oracle then enhances the amplitude of the state that represent the correct location. By repeatedly applying the Grover operator, the probability amplitude of the correct position is amplified with each iteration, making it more likely to be measured upon observation.
In practical terms (if real quantum computing achieves enough capacity), this quantum-enhanced localization approach could be particularly beneficial in environments with high location ambiguity or large search spaces, such as a warehouse with many aisles or complex navigation zones. In principle, leveraging quantum computing’s ability to quickly process great amounts of data (exponentially faster than classical computers), robots and autonomous vehicles can quickly determine their precise location and orientation, even in situations where GPS signals are weak or unavailable.
@@
That is why in this paper we propose using Grover’s quantum search oracle to perform a double localization on a 2D map: one for the horizontal axis and another for the vertical axis, as shown in Figure 9
Figure 9: Block diagram of the proposed approach
@@
We left aside other current approaches for pattern matching, like the use of Parametric Probabilistic Quantum Memory ((Sousa et al., 2020)) (PQM, a data structure that computes the Hamming distance from a binary input to all binary patterns stored in superposition on the memory), as they require the substring to be searched to be of the same length as the strings in the stored data.
@@ 3.1 Algorithm
Figure 10: The perception costmap (left) has 3 horizontal elements, and 3 vertical elements. The map (the previously generated costmap, in the right) is a 4x4 grid of cells (each cell has two qubits)
@@
Figure 11: Considering (0,0) as the upper left cell, for coordinates (1,1) if the map’s cell values (in black) match the values for the horizontal (blue) and vertical (red), the Oracle’s output should be ’1’, and ’0’ otherwise. In other words: the algorithm should look for [11,10,11] horizontally, and [11,10,10] in vertical.
The code generated in this paper ((Antero, 2024)) can be configured for any map size, any number of qubits per cell, and any row or column pattern. In this case (with a 4x4 map with 2 qubits per cell and horizontal and vertical patterns with size 3 (using 2 qubits per cell), the Oracle is automatically generated. Such an oracle (in this case) has a possible search space of: [(1,1), (1,2), (2,1), (2,2)], and would return ’1’ in the output qubit for position: (1,1).
@@ 4 Experimental Setup and Validation
In the preceding sections, we introduced a novel application of quantum computing to the problem of robot localization, proposing a quantum algorithm designed to efficiently solve the localization problem. To validate the feasibility and effectiveness of this quantum approach, this section presents a comprehensive experimental setup and validation framework. We outline the methodology employed to evaluate the performance of our quantum algorithm in estimating a robot’s position and orientation within a given environment and present the results obtained from a series of simulations and experiments. These experiments aim to assess the accuracy, robustness, and scalability of our quantum algorithm, and demonstrate its potential advantages over classical methods for robot localization. The remainder of this section details the experimental design, performance metrics, and results, offering a thorough analysis of the proposed quantum approach to robot localization and its potential applications in robotics and autonomous systems.
@@
The first module is an auxiliary configuration tool developed using Pygame ((pyg, 2024)) (a Python library designed for creating video games and multimedia applications). This tool enables users to define and customize the parameters of the localization problem, including the 2D map, robot properties, and algorithmic settings. The output of the tool is a standardized input file that serves as the basis for the second module.
@@
  • Create a map with a size $n\times n$, being $n\geq 2$
@@
Once launched, the application shows a simple user interface where, using arrow keys, it is possible to move the robot to any cell in the map.
@@
Using this user interface, different setups (described in the following subsections) have been prepared and tested. These setups cover different situations: from small grids that were simulated locally in a PC, to more complex grids that had to be simulated in real quantum computers.
On the other hand, within the scope of this paper, we have also implemented the algorithm described in 3. This main module, built using the Qiskit quantum development environment, implements the quantum algorithm for robot localization in a 2D map.
@@
By decoupling the problem configuration from the implementation of the quantum algorithm, our framework provides a versatile and user-friendly environment for testing and evaluating the performance of the proposed quantum algorithm for robot localization.
In any case, it should be noted that quantum computing, as a technology, is still in its infancy. It is clear that different engineering challenges in the real quantum computer disable (in a certain sense) any advantage that the proposed algorithm may present. In other words: notwithstanding the theoretical validity of the proposed quantum approach, the practical implementation of the corresponding quantum circuit is subject to the engineering limitations of current quantum computing technology.
@@
In real quantum computers, we currently face four main challenges:
  • Stability of the Qubit state: The stability of a qubit’s state typically lasts only a few hundred microseconds, which limits the duration of computations.
@@
  • Limited qubit availability: The number of physical qubits available in these devices remains constrained, which poses a significant bottleneck for scalability.
Limited qubit availability impacts on the complexity of the algorithms we can use, but the other three issues directly impact the error probability of the final result. As indicated in ((Johnstun and Huele, 2021)), it is important to understand the effect of these issues and to have some estimation of the final error.
In this paper, we use data provided by the manufacturer of the real quantum computer (IBM in our case), to calculate the total error probability expected during the run of a given quantum circuit. Using the methodology (and code) proposed in ((Aseginolaza et al., 2024)), for each setup we estimate how far we are from the limits of current quantum computers (for using their code in this paper, the code has been updated to the latest version of Qiskit).
In the next subsections, we describe the experimental setups and results obtained using the software developed in this paper, which demonstrates the effectiveness of our quantum approach to robot localization (and the limits of current physical quantum computers).
4.1 Small grid with simple pattern match
As a first test for validation, a very simple setup has been configured: it is a 2x3 qubit grid that uses a single qubit per cell to encode the costmap.
The Pygame based interface has been used to configure such a grid, and to define the robot’s position in a certain cell in the grid.
An obstacle has been placed in an arbitrary location, as seen in Figure 13. Now, the goal is to find the location in which a robot would perceive such an obstacle.
Map qubits: $\begin{bmatrix}0&1&0\\ 0&0&0\\ \end{bmatrix}\$
@@ 4.1 Small grid with simple pattern match
Figure 13: Simple setup with a 2x3 grid
The proposed double (horizontal and vertical) Grover search with a pattern $[1]$ in a $2\times 3$ grid, has to deal with a search space of 6 qubits. For this setup, our algorithm requires us to use:
  • The Oracle has to deal with 6 possible positions (2x3), with the expectation of a single correct solution (N=6, M=1).
  • This requires 3 qubits for the search space (to encode 6 positions)
  • The suggested number of repetitions is: $R=\pi/4\sqrt{N/M}=2$
  • The total number of required qubits is 13
  • Size of the quantum circuit is 163 elements (with a depth of 89)
In order to validate the effectiveness of the algorithm, the Qiskit code has been tested on different systems (simulated and real):
  • Using the IonQ Quantum Cloud ((ion, 2024)), where it is possible to simulate circuits as if they were running on IonQ hardware, using a special noise model simulation.
  • Using the Bluequbit Cloud platform ((blueqbit)), a platform specialized in the simulation of quantum circuits .
  • Using a simulation based on the ’qasm’ simulator, but where noise and error rates of a real quantum computer have been introduced. In this test, the error rates of IBM’s "ibm brisbane" quantum processing unit weree used (a 127qubit unit, with a 2Q error (best) of 3.28e-3).
  • Using a real quantum computer from IBM ((ibm, 2024b)). Using IBM’s SDK it is possible to send a quantum circuit to a real IBM quantum computer. In this test, a real "ibm brisbane" quantum processing unit was also used for the execution of the algorithm (the code was sent remotely to be executed in a job queue).
Grover’s algorithm identified the correct solution: row 1 and column 2, identified by position ’0’ or ’001’ in binary (count starts from zero on the upper left corner of the grid). The probability of the correct solution was amplified in the simulated environment and in the real environment.
However, since these operations are performed in the realm of quantum mechanics, governed by the principles of superposition, entanglement, and interference, there is the possibility that the results found vary from the calculations in the ideal environment.
This could be computed due to the existence of noise, since noise refers to random errors that occur during computation.
These errors cause the qubits to lose their quantum properties and behave classically. Consequently, as a direct result, actual computations diversify from the calculated probability and from simulated equations.
Furthermore, real quantum computers operate differently from ideal simulations because of their susceptibility to noise.
In the presence of noise, the results of the computation may tend to deviate from the results in the ideal simulation, consequently causing discrepancies between the results and expectations of behavior from the quantum system.
The histograms obtained are shown in Figure 14:
(a) Simulated on Bluequbit
(b) Simulated on IonQ
(c) Simulated using "ibm brisbane" unit’s noise model
(d) Run on a real IBM Quantum computer
Figure 14: Histograms of the same quantum circuit obtained from simulated and real quantum computers
It is clear that the noise in the real quantum computer masks (in a certain sense) any advantage that the proposed algorithm may present. In the tests performed, although most of the time the results are correct, sometimes the algorithm tends to fail on the real computer. In the "ibm brisbane" quantum unit that has been used, the lowest two-qubit error rate (as indicated by IBM) is 3.28e-3. Using this information, it is possible to have a rough estimate of the probability that the algorithm ran without any error (using the ’depth’ as the number of layers that work one after the other):
$\begin{split}P(\text{at least 1 gate fails})&=1-P(\text{all gates succeed})\\ &=1-P(\text{one gate succeeds})^{\text{number of gates}}\\ &=1-(1-P(\text{gate error}))^{\text{aprox. depth of circuit}}\\ &=1-(1-3.28e-3)^{89}\\ &\approx 25\%\end{split}$ (4)
This value is aligned with how the expected error is estimated according to (Aseginolaza et al., 2024) whose code has been integrated in the code of this paper and estimates that for this 2x3 grid we get:
  • Probability of error due to noisy gates = 28%
  • Probability due to measurement errors = 18%
4.2 Bigger grid, with simple pattern match
To perform a more useful validation, it is necessary to perform a simulation that is computationally more demanding, so in this setup a 5x5 grid will be used. The setup considers a 1-qubit pattern search (both horizontally and vertically), and the calculated costmap in each cell has been configured to use a single qubit. Once again, the Pygame-based interface has been used. The robot has been placed in the central cell, as seen in Figure 15.
Map qubits: $\begin{bmatrix}0&0&0&0&0\\ 0&0&0&0&0\\ 0&0&1&0&0\\ 0&0&0&0&0\\ 0&0&0&0&0\\ \end{bmatrix}\$
Horizontal search qubits: $\begin{bmatrix}1\end{bmatrix}\\ \$
Vertical search qubits: $\begin{bmatrix}1\end{bmatrix}\$
Figure 15: Simple setup with a 5x5 grid
The proposed double Grover search with a pattern $[1]$ has to deal with a search space of $55x$ 5. In order to evaluate the quantum circuit, the proposed algorithm needs:
  • 5 qubits for the search space (to encode positions 0 to 24)
  • The Oracle has to deal with 4 possible positions (5x5), with an expectation of a single correct solution (N=25, M=1)
@@ 4.2 Bigger grid, with simple pattern match
  • The total number of required qubits is 34
  • Circuit depth is 707
@@
  • using the IonQ cloud for a simulated result
  • using Bluequbit for a simulated result
@@
The real IBM was the "ibm kyiv" machine, a 127qubit machine with a "2Q Error" rate (lowest two-qubit error rate) of 3.81e-3.
(a) Simulation in IonQ
(b) Simulation in Bluequbit
(c) Real quantum computer in ’bm kyiv’ quantum unit. Noise in the system hides de amplification
@@
Figure 16: Histograms of the same quantum circuit obtained from simulated and real quantum computers
For the "ibm kyiv" machine, the probability to run the algorithm without any error is:
$\begin{split}P(\text{at least 1 gate fails})&=1-(1-3.81e-3)^{707}\\ &\approx 93\%\end{split}$ (5)
Once again, this value is aligned with the expected error according to (Aseginolaza et al., 2024) whose code has been integrated in the code of this paper and estimates that for this grid we would get:
  • Probability of error due to noisy gates ${\approx 99\%}$
  • Probability due to measurement errors = 73%
4.3 Grid with a pattern match with two qubits